I need to solve the following system of equations
x+6y=9 (1)
x=7-6y (2)
What is the solution to the system?


Thank you

Answer by mananth(16946)

x+6y=9 (1)
x=7-6y (2)
add 6y to both sides
x+6y=7... (2)
compare (1) & (2)
we get 9=7
...
There is no solution to this system